OpenStudy (anonymous):

What is the length of a in this triangle? 2 12 35.5 8

OpenStudy (anonymous):



OpenStudy (anonymous):

do you know Pythagorean theorem?

OpenStudy (anonymous):

Yes A^2+B^2+C^2

OpenStudy (anonymous):

what? a^2 + b^2 + c^2? no,a^2 + b^2 = c^2 not +

OpenStudy (anonymous):



OpenStudy (anonymous):

I know everything up to A^2, then I get confused

OpenStudy (wolf1728):

In this case 37² -35² = a²

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

yes but if you still get confused maybe the proof of Pythagorean theorem would help you.

OpenStudy (anonymous):

So then youo get 2,594= A^2 ?

OpenStudy (wolf1728):

Bookreader what is 37² ? What is 35²

OpenStudy (anonymous):

isnt 37X37 or am I having a brain fart?

OpenStudy (wolf1728):

Yes it is 37*37 and what is it?

OpenStudy (anonymous):

1,369

OpenStudy (wolf1728):

and what is 35²?

OpenStudy (anonymous):

1,225

OpenStudy (wolf1728):

And subtracting 1,225 from 1,369 we get ?

OpenStudy (anonymous):

144

OpenStudy (anonymous):

and 144 is 12^2 :)

OpenStudy (wolf1728):

and 144 is a squared so 'a' equals?

OpenStudy (anonymous):

12

OpenStudy (wolf1728):

Very good !!!

OpenStudy (anonymous):

Thank you so much!

OpenStudy (wolf1728):

u r welcome!!!